OpenStudy (anonymous):

PLEASE HELP!!(: For events A and B, P(A)=4/13 and P(B)=1/4 . Also, P(A and B)=1/15 . Are A and B independent? Please explain.

OpenStudy (anonymous):

if they are independent then \[P(A\cap B)=P(A)\times P(B)\]

OpenStudy (anonymous):

so check to see if, when you multiply, you get \(\frac{1}{15}\)

OpenStudy (anonymous):

looks to me like you do not, since \(\frac{4}{13}\times \frac{1}{4}=\frac{1}{13}\neq \frac{1}{15}\)

OpenStudy (anonymous):

@satellite73 so then what does that conclude?

OpenStudy (anonymous):

@apoorvk this one

OpenStudy (apoorvk):

yeah satellite is right there. A and B are not independent events. now, does "and" there mean union or intersection? and, first, the vein diagram!